Find the solution to the system of equations.
2x+4y+2z=16
3x+2y+2z=13
-4x+5y+3z=15

Answer by Edwin McCravy(20054)

2x+4y+2z=16
3x+2y+2z=13
-4x+5y+3z=15
There are many ways to solve that system. The solution is
(x,y,z) = (1,2,3)
Post again telling which method you are studying and we
can help you with solving it, but there are many
different ways to solve it which are studied. For
us to just guess one of the many ways and solve it that
way would likely not be the way you are studying.